40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Sep 10 02:53:32 2014, Jenne, Update, LSC, Locking activities - nothing new :( LocklossZoom.png
    Reply  Wed Sep 10 02:59:42 2014, ericq, Update, LSC, Locking activities - nothing new :( 
Message ID: 10485     Entry time: Wed Sep 10 02:53:32 2014     Reply to this: 10486
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Locking activities - nothing new :( 

[Jenne, EricQ]

No major progress today.  

I fixed a bug in my lockloss script that was asking it to start gathering data just after the lockloss, rather than some seconds beforehand.  Ooops.  Anyhow, with this handy-dandy plotting, I still don't know why we are losing lock when we have PRMI on REFL33, CARM on sqrtInvTrans, and DARM on AS55.  I don't see any oscillations, just the arm power drops off, and a moment later the POP power drops.  

For example, here is one of the best states we got to tonight.  Data for this is in ..../scripts/LSC/LocklossData/1094369700 .  You can re-create the plot by going to ..../scripts/LSC/LocklossData/  and doing ./PlotLockloss.py 1094369700 .  We had set the triggers for the trans PD/QPD such that we were using the QPD transmission signals the whole time (above trans of 0.2).  We saw that the noise at high frequency during low transmission powers for sqrtInvTrans as an error signal was higher using the QPDs than with the Thorlabs PDs, but that both cases are below the noise for ALS.  The arm powers were pretty steady above 3 for the last bit of this lock stretch.  I lost lock while trying to transition DARM over to AS55Q.  CARM was on sqrtInvTrans(QPDs), PRMI on REFL33 I&Q as usual.

LocklossZoom.png

 


Other things from this evening:

* When I was starting, I saw that when I locked the PRMI, the PRM was oscillating in pitch. Oscillation only happened when PRM pitch oplev was on.  I'm not sure what could have changed to make the oplev loop unstable, but the gain was 7.0, and now I have left it at 5.0.

* I recentered the PRM and ITMY oplevs.

* Plugged in the Yend PDH error monitor and pzt output monitors, since I forgot them last week.  Hopefully this will allow the Yend SLOW servo to work, and keep us away from the limits of the PZT range.

ELOG V3.1.3-